OpenStudy (anonymous):

Prove:

mathslover (mathslover):

Hello @Yaho021 , please post your problem so that we can help you ... If you have posted this question by fault or you have not anything to ask , then please close this question .

OpenStudy (anonymous):

\[e ^{i \pi}+ 1 = 0\]

mathslover (mathslover):

OK can you tell me that what is \(\large{e^{i \pi}}\) ?

mathslover (mathslover):

@yaho021 Please reply to my question.

OpenStudy (unklerhaukus):

\[e^{inx}=\cos(nx)+i\sin(nx)\] \[e^{iπ}=\cos(π)+i\sin(π)=-1+0i\]

mathslover (mathslover):

Oh you disclosed the answer , I thought that the asker will tell it by his/her own.